OpenStudy (anonymous):

Find the x-intercept and the y-intercept of the graph of the equation. 2x + y = -8

OpenStudy (mathmale):

Dear Bobie: Think of the x-intercept as the point on the x-axis where the line 2x+y=-8 crosses that axis. At that point, y=0. So: let y = 0 in 2x + y = -8. Solve the resulting value for x.

OpenStudy (anonymous):

As mathmale have said for x-intercept put y=0 and for y-intercept put x=0 in the equation. @mathmale good work bro!

OpenStudy (anonymous):

Try it for yourself & post the answers we'll check them !

OpenStudy (anonymous):

let y=0 therefore 2x + 0 =-8 x=-4 so is that the x intercept?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome